18 - 5r - 5 = -17 solve for r
Answer by rameens_99(10)

18-5r-5= -17 (combine like terms that are on the SAME side of the equal sign)
13-5r= -17 (now subtract 13 from both sides)
-5r= -30 (divide -5 by -30)
_________________
So r= 6





You're Welcome (-: